OpenStudy (anonymous):

determine whether the lines are parallel, perpendicular or neither. 1. One line passes through the points (-1, 4) and (2, 6); another line passes through the points (2, -3) and (8, 1). 2. One line passes through the points (4, -3) and (-8, 0); another line passes through the points (-1, -1) and (-2, 6). 3. One line passes through the points (-3, 14) and (1, -2); another line passes through the points (0, -3) and (-2, 5).

OpenStudy (anonymous):

Ok to do this problem you would first use slope intercept form to write the equation for each line

OpenStudy (acxbox22):

do you know how to calculate slope?

OpenStudy (anonymous):

no i dont

OpenStudy (acxbox22):

slope is the measure of the constant change in a linear equation you can find slope when you are given 2 points (x1,y1) and ((x2,y2) you calculate slope by using formula y2-y1 ----- x2-x1

OpenStudy (anonymous):

Gosh why do you guys have to talk so nerdy lol

OpenStudy (anonymous):

All slope is is change in your y coordinate/change in x coordinate

OpenStudy (acxbox22):

find slopes of each 2 lines in each question if the slopes are same then they are parallel if the slopes are opposite reciprocals (ex. slope=1 slope=-1/2) they are perpendicular.

OpenStudy (anonymous):

So if you plot your points on a graph for the first one

OpenStudy (acxbox22):

you have to talk nerdy the first time cause she doesnt know what slope is at all

OpenStudy (anonymous):

you see that y varies from 4 to 6. So chance in y is 2

OpenStudy (anonymous):

x varies from -1->2 so the change in x is 3

OpenStudy (anonymous):

therefore the slope is 2/3

OpenStudy (anonymous):

Ok and then use the equation y-y1=m(x-x1)

OpenStudy (anonymous):

here you replace (x1,y1) with any one of the points. In this case I will use (-1,4)

OpenStudy (anonymous):

and the calculated slope of 2/3 in place of m

OpenStudy (anonymous):

arriving at y-4=2/3(x+1)

OpenStudy (anonymous):

Let me know if you need any more help